OpenStudy (anonymous):

lim x->0 sin3x/5x^3-4x.

OpenStudy (shamim):

\[\lim_{x \rightarrow 0}\frac{ \sin 3x }{ 5x ^{2}-4 }=\frac{ \lim_{x \rightarrow 0}3x \frac{ \sin 3x }{ 3x } }{ 5x ^{3}-4x }\]

OpenStudy (shamim):

did u get it

OpenStudy (shamim):

\[=\frac{ 3x \lim_{x \rightarrow 0}\frac{ \sin 3x }{ 3x } }{ 5x ^{3}-4x }\]

OpenStudy (shamim):

u know\[\lim_{x \rightarrow 0}\frac{ \sin 3x }{ 3x }=1\]

OpenStudy (shamim):

now try to do the rest

OpenStudy (anonymous):

so would you want to factor out the x?

OpenStudy (shamim):

ya

OpenStudy (anonymous):

got it thanks

OpenStudy (shamim):

\[\lim_{x \rightarrow 0}\frac{ 3x }{ 5x ^{3}-4x }\]

OpenStudy (shamim):

welcome

OpenStudy (shamim):

medal????????????????

OpenStudy (anonymous):

i dont know how to do that or what means.. I am sorry I just found this site

OpenStudy (shamim):

just click on best answer

OpenStudy (anonymous):

Hey did you by any chance see my other question?

OpenStudy (shamim):

ya